OpenStudy (anonymous):

A particle is moving on the x-axis has displacement x(t)=2t^3+3t^2-36t+40 after an elapsed time of t seconds. a) Find the velocity of the particle at time t. b) Find the acceleration of the particle at time t.

OpenStudy (anonymous):

velo=v=dx/dt=6t^2+6t-36 acce=dv/dt=12t+6